Four coins are tossed. Let M be the variable representing the number of heads that occur. Find the values of the rand om variable M and its corresponding probabilities

1
Expert's answer
2022-02-28T13:25:44-0500

If 4 coins are tossed, then there can be from 0 to 4 heads. M has a binomial distribution with n = 4, p = 0.5, so

"P(M=0)={4 \choose 0}*({\frac 1 2})^4={\frac 1 {16}}"

"P(M=1)={4 \choose 1}*({\frac 1 2})^4={\frac 4 {16}}={\frac 1 4}"

"P(M=2)={4 \choose 2}*({\frac 1 2})^4={\frac 6 {16}}={\frac 3 8}"

"P(M=3)={4 \choose 3}*({\frac 1 2})^4={\frac 4 {16}}={\frac 1 4}"

"P(M=4)={4 \choose 4}*({\frac 1 2})^4={\frac 1 {16}}"